OpenStudy (anonymous):

Please please help c: Thanks What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10 (−4, −2) and (2, 5) (−2, −4) and (2, 5) (−2, −4) and (5, 2) No Real Solutions

OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + 10x + 11 y = x2 + x − 7 (−2, −5) (2, −5) (−2, 5) (2, 5)

OpenStudy (anonymous):

Try Elimination method to solve these equations

OpenStudy (debbieg):

You have 2 different expressions for y. So if both of those things are =y, then they are also equal one another. So set the expressions =, then solve for x. The x^2 terms will conveniently go away, and you will have a simple linear equation in x to solve. Once you have solved that, you will know the value of x. Then you can just substitute that value back into either of the equations to find y.